# On the ring of cooperations for Real Hermitian K-theory

### v1

@@ 6.2. Some modules over $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$
The big flag $\mathcal{F}_{0,0,0}$ is isomorphic to the coweight $cw\equiv 0\,(4)$ piece of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ (see Fig. 6).
@@ 6.3. Comparison with $\mathbb{C}$
There is a well-studied relationship between $\text{SH}(\mathbb{C})$ and $\text{SH}(\mathbb{R})$ [BI22]. In [Hil11], Hill developed the $\rho$-Bockstein spectral sequence, an algebraic spectral sequence which allows for the computation of $\mathbb{R}$-motivic information from $\mathbb{C}$-motivic information. The observation that $\mathbb{M}_{2}^{\mathbb{C}}=\mathbb{M}_{2}^{\mathbb{R}}/\rho$ lifts to the dual Steenrod algebra, giving $\mathcal{A}^{\vee}_{\mathbb{C}}=\mathcal{A}^{\vee}_{\mathbb{R}}/\rho$. Filtering the cobar complex which computes $\text{Ext}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ by powers of $\rho$, we obtain the $\rho$-Bockstein spectral sequence
$E_{1}=\text{Ext}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})[\rho]\implies\text{Ext}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}}).$
Note that this spectral sequence converges to the $E_{2}$-page of the $\mathbb{R}$-motivic mASS($\mathbb{S}$). One can also filter the cobar complex which computes $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ by powers of $\rho$, giving a $\rho$-Bockstein spectral sequence
$E_{1}=\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})[\rho]\implies\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}}).$
[…]
There is also an extension of scalars functor $-\otimes\mathbb{C}:\text{SH}(\mathbb{R})\to\text{SH}(\mathbb{C})$. For $X\in\text{SH}(\mathbb{R})$, we denote its image under extensions of scalars by $X^{\mathbb{C}}$. This functor has the property $\rho\otimes\mathbb{C}=0$ (modulo 2) and Eilenberg-Mac Lane spectra to are sent to Eilenberg-Mac Lane spectra. In particular, this gives maps
@@
In the 2-complete setting, one may lift $\rho$ to the sphere, giving an element $\rho\in\pi_{-1,-1}^{\mathbb{R}}\mathbb{S}$. This gives a cofiber sequence
[…]
@@
[…]
@@
$\textup{Ext}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\textup{H}_{**}(X/\rho))\xrightarrow{\cong}\textup{Ext}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\textup{H}_{**}(X^{\mathbb{C}})).$
@@
One immediate consequence of Proposition 6.6 is that $\text{Ext}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$ is a module over $\text{Ext}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2})$. This module structure is simple to describe. By the $\rho$-Bockstein, we can represent classes in $\text{Ext}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ in the form $\rho^{k}x$ for some $x\in\text{Ext}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$. If $y\in\text{Ext}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$, then $\rho^{k}x\cdot y$ is only nonzero when $k=0$, and in this case we have $x\cdot y=xy$.
@@
The map from Proposition 6.6 gives an isomorphism of $E_{2}$-pages of Adams spectral sequences. ∎
Note that $\textup{kq}^{\mathbb{C}}$ is the spectrum in $\text{SH}(\mathbb{C})$ representing very effective Hermitian K-theory over $\mathbb{C}$, which we will abusively also denote by kq. Since $\textup{H}_{**}(\textup{kq})\cong(\mathcal{A}_{\mathbb{R}}//\mathcal{A}(1))_{\mathbb{R}}^{\vee}$ is $\rho$-torsion free, it follows from Corollary 6.7 that $\pi_{**}^{\mathbb{R}}(\textup{kq}/\rho)\cong\pi_{**}^{\mathbb{C}}(\textup{kq})$. The Künneth spectral sequence for $\textup{H}_{**}(\textup{kq}\otimes\textup{kq})$ collapses by Proposition 2.12, so we also have
@@
The $\rho$-periodic $E_{1}$-page of the real kq-resolution is isomorphic to the $E_{1}$-page of the complex kq-resolution.
We can also prove analogous statements using the Brown–Gitler comodules. First, note that we have a diagram in Ext coming from the long exact sequences induced by the cofiber sequence
[…]
This gives an isomorphism in Ext by the same argument as before:
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}}/\rho)\cong\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}}).$
By construction, the Brown–Gitler comodules $B_{0}^{\mathbb{R}}(k)$ are $\rho$-torsion free since they are subcomodules of the dual Steenrod algebra. This extends naturally to tensor factors leading us to the following result:
@@
$\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k)^{\otimes i}/\rho)\cong\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k)^{\otimes i}).$
These isomorphisms will be helpful for deducing hidden extensions on the $E_{\infty}$-page of the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(k))$ by comparing with the computations of [CQ21] recalled in Section 5.
@@
obtained from the Betti realization functor of Remark 2.6 by remembering the $C_{2}$-action given by complex conjugation. This functor has been well studied, see for example [HO16], and exhibits a close relationship between $\mathbb{R}$-motivic and $C_{2}$-equivariant homotopy theory. It was shown by [Kon23] that
$\text{Be}^{C_{2}}(\textup{kq}_{2,\eta}^{\wedge})\simeq(\text{ko}_{C_{2}})_{2,\eta}^{\wedge}.$
@@
$\text{Ext}_{\mathcal{A}^{\vee}_{C_{2}}}\cong\text{Ext}_{\mathcal{A}_{\mathbb{R}}^{\vee}}\oplus\text{Ext}_{NC}.$
[…]
@@ 7. The algebraic Atiyah–Hirzebruch spectral sequence
In this section, we compute the group $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes i})$ by a series of algebraic Atiyah–Hirzebruch spectral sequences. Motivated by the presentation of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ given in [GHIR20], we organize these computations by coweight modulo 4. Consequently, our results are best displayed in multiple charts which should be read simultaneously. Throughout, we use the notation of Section 6.
7.1. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$
Now, we will compute $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$ by the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$. This spectral sequence takes the form
$E_{1}^{s,f,w,a}=\text{Ext}^{s,f,w}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})\otimes_{\mathbb{M}_{2}^{\mathbb{R}}}\mathbb{M}_{2}^{\mathbb{R}}\{[1],[\overline{\xi}_{1}],[\overline{\tau}_{1}]\}\implies\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}^{s,f,w}(B_{0}^{\mathbb{R}}(1)).$
@@ 7.1. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$
and the $d_{2}$-differential is only nonzero on Atiyah–Hirzebruch filtration 3, where
@@
Note that in terms of coweight, both of these differentials are determined by multiplication by an element in $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ of coweight 0. Given the complicated nature of this spectral sequence, we only present charts where the differential is nonzero and only include the relevant Atiyah–Hirzebruch filtrations. For example, since the $d_{1}$-differential is trivial on Atiyah–Hirzebruch filtration 0, we omit this filtration from the charts for the $E_{1}$-page.
Fig. 10 depicts the $E_{1}$-page with coweight $cw\equiv 0\,(4)$ in Atiyah–Hirzebruch filtrations 2 and 3. Fig. 11 depicts the $E_{1}$-page with coweight $cw\equiv 1\,(4)$ in Atyiah-Hirzebruch filtrations 2 and 3. Fig. 12 depicts the $E_{1}$-page with coweight $cw\equiv 2\,(4)$ in Atiyah–Hirzebruch filtrations 2 and 3. Fig. 13 depicts the $E_{2}$-page with coweight $cw\equiv 0\,(4)$ in Atiyah–Hirzebruch filtrations 0 and 2. Fig. 14 depicts the $E_{2}$-page with coweight $cw\equiv 1\,(4)$ in Atiyah–Hirzebruch filtrations 0 and 2. Fig. 15 depicts the $E_{2}$-page with coweight $cw\equiv 2\,(4)$ in Atiyah–Hirzebruch filtrations 0 and 2. Differentials are blue, linear with respect to the underlying $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure, and preserve motivic weight.
Figure 10. The $E_{1}$-page of the aAHSS($B_{0}^{\mathbb{R}}(1)$) with $cw\equiv 0\,(4)$ in Atiyah–Hirzebruch filtration 2 and 3.
Figure 11. The $E_{1}$-page of the aAHSS($B_{0}^{\mathbb{R}}(1)$) with $cw\equiv 1\,(4)$ in Atiyah–Hirzebruch filtration 2 and 3.
Figure 12. The $E_{1}$-page of the aAHSS($B_{0}^{\mathbb{R}}(1)$) with $cw\equiv 2\,(4)$ in Atiyah–Hirzebruch filtration 2 and 3.
Figure 13. The $E_{2}$-page of the aAHSS($B_{0}^{\mathbb{R}}(1)$) with $cw\equiv 0\,(4)$ in Atiyah–Hirzebruch filtration 0 and 2.
Figure 14. The $E_{2}$-page of the aAHSS($B_{0}^{\mathbb{R}}(1)$) with $cw\equiv 1\,(4)$ in Atiyah–Hirzebruch filtration 0 and 2.
Figure 15. The $E_{2}$-page of the aAHSS($B_{0}^{\mathbb{R}}(1)$) with $cw\equiv 2\,(4)$ in Atiyah–Hirzebruch filtration 0 and 2.
@@
This differential is witnessed by the Massey product $\langle\rho,h_{0},h_{1}\rangle$. This Massey product is shown to be equal to $\tau h_{1}$ in [GHIR20, Thm 8.1] with zero indeterminacy. ∎
In fact, since the differentials are linear over $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ this determines all $d_{3}$-differentials. For example, we have
@@
It is important to note that since $\tau h_{1}$ has coweight 1 in $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$, and since each Atiyah–Hirzebruch filtration is zero in coweight 3, the $d_{3}$-differential is trivial on the coweight $cw\equiv 2\,(4)$ portion of Atiyah–Hirzebruch filtration 3. This coweight jump requires us to pay close attention to the $d_{3}$-differential. For example, in the coweight $cw\equiv 0\,(4)$ piece of Atiyah–Hirzebruch filtration 3, there is an infinite $\rho$-tower stemming from $\rho[3]$. However, while
@@
Fig. 16 depicts the $E_{3}$-page where the coweight of Atiyah–Hirzebruch filtration 3 is $cw\equiv 0\,(4)$ and the coweight of Atiyah–Hirzebruch filtration 0 is $cw\equiv 1\,(4)$. Fig. 17 depicts the $E_{3}$-page where the coweight of Atiyah–Hirzebruch filtration 3 is $cw\equiv 1\,(4)$ and the coweight of Atiyah–Hirzebruch filtration 0 is $cw\equiv 2\,(4)$.
Figure 16. The $E_{3}$-page of the aAHSS($B_{0}^{\mathbb{R}}(1)$) with $cw\equiv 0\,(4)$ in Atiyah–Hirzebruch filtration 3 and $cw\equiv 1\,(4)$ in Atiyah–Hirzebruch filtration 0.
Figure 17. The $E_{3}$-page of the aAHSS($B_{0}^{\mathbb{R}}(1)$) with $cw\equiv 1\,(4)$ in Atiyah–Hirzebruch filtration 3 and $cw\equiv 2\,(4)$ in Atiyah–Hirzebruch filtration 0.
Recall from Proposition 4.6 that $E_{4}=E_{\infty}$ for Atiyah–Hirzebruch filtration reasons. Hidden extensions can be realized using complex Betti realization, linearity over $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$, base change $\otimes\mathbb{C}:\text{SH}(\mathbb{C})\to\text{SH}(\mathbb{R})$, and the isomorphism from Proposition 6.9:
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)/\rho)\cong\text{Ext}_{\mathcal{A}(1)_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)).$
The following result will be useful when depicting the $E_{\infty}$-page in charts.

### v2

@@ 6.1. Background
The big flag $\mathcal{F}_{0,0,0}$ is isomorphic to the coweight $cw\equiv 0\,(4)$ piece of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ (see Figure 6.1).
Remark 6.5.
The big flag $\mathcal{F}_{4n,n,2n}$ is a submodule of the coweight $cw\equiv 0\,(4)$ portion of the $\text{E}_{2}$-page of the $\textbf{mASS}^{\mathbb{R}}(\text{kq}^{\langle n\rangle})$, where $\text{kq}^{\langle n\rangle}$ denotes the $n^{th}$ Adams cover.
@@ 6.3. Comparison with $\mathbb{C}$
There is a well-studied relationship between $\text{SH}(\mathbb{C})$ and $\text{SH}(\mathbb{R})$ [12]. In [26], Hill developed the $\rho$-Bockstein spectral sequence, an algebraic spectral sequence which allows for the computation of $\mathbb{R}$-motivic information from $\mathbb{C}$-motivic information. The observation that $\mathbb{M}_{2}^{\mathbb{C}}=\mathbb{M}_{2}^{\mathbb{R}}/\rho$ lifts to the dual Steenrod algebra, giving $\mathcal{A}^{\vee}_{\mathbb{C}}=\mathcal{A}^{\vee}_{\mathbb{R}}/\rho$. Filtering the cobar complex which computes $\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ by powers of $\rho$, we obtain the $\rho$-Bockstein spectral sequence
$\textup{E}_{1}=\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})[\rho]\implies\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}}).$
This spectral sequence converges to the $\textup{E}_{2}$-page of the $\textbf{mASS}^{\mathbb{R}}(\mathbb{S})$. One can also filter the cobar complex which computes $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ by powers of $\rho$, giving a $\rho$-Bockstein spectral sequence
$\textup{E}_{1}=\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})[\rho]\implies\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}}).$
This spectral sequence converges to the $\textup{E}_{2}$-page of the $\textbf{mASS}^{\mathbb{R}}(\textup{kq})$.
There is also an extension of scalars functor $\otimes\mathbb{C}:\text{SH}(\mathbb{R})\to\text{SH}(\mathbb{C})$. For $X\in\text{SH}(\mathbb{R})$, we denote its image under extensions of scalars by $X^{\mathbb{C}}$. This functor has the property $\rho\otimes\mathbb{C}=0$ (modulo 2) and Eilenberg-Mac Lane spectra to are sent to Eilenberg-Mac Lane spectra. In particular, this gives maps
@@
In the 2-complete setting, one may lift $\rho$ to the sphere, giving an element $\rho\in\pi_{-1,-1}^{\mathbb{R}}(\mathbb{S})$. This gives a cofiber sequence
$\Sigma^{-1,-1}\mathbb{S}\xrightarrow{\rho}\mathbb{S}\to\mathbb{S}/\rho.$
@@
$\cdots$$\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$$\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$$\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}}/\rho)$$\cdots$$\cdots$$\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$$\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$$\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})\oplus\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$$\cdots$$\scriptstyle{\rho}$$\scriptstyle{0}$
@@
$\textup{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\textup{H}_{**}(X/\rho))\xrightarrow{\cong}\textup{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\textup{H}_{**}(X^{\mathbb{C}})).$
@@
One immediate consequence of Proposition 6.2 is that $\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$ is a module over $\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2})$. This module structure is simple to describe. By the $\rho$-Bockstein, we can represent classes in $\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ in the form $\rho^{k}x$ for some $x\in\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$. If $y\in\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$, then $\rho^{k}x\cdot y$ is only nonzero when $k=0$, and in this case we have $x\cdot y=xy$.
@@
The map from Proposition 6.2 gives an isomorphism of $\textup{E}_{2}$-pages of Adams spectral sequences. ∎
Note that $\textup{kq}^{\mathbb{C}}$ is the spectrum in $\text{SH}(\mathbb{C})$ representing very effective Hermitian K-theory over $\mathbb{C}$, which we will abusively also denote by kq. Since $\textup{H}_{**}(\textup{kq})\cong(\mathcal{A}\operatorname{/\!/}\mathcal{A}(1))_{\mathbb{R}}^{\vee}$ is $\rho$-torsion free, it follows from Corollary 6.3 that $\pi_{**}^{\mathbb{R}}(\textup{kq}/\rho)\cong\pi_{**}^{\mathbb{C}}(\textup{kq})$. The Künneth spectral sequence for $\textup{H}_{**}(\textup{kq}\otimes\textup{kq})$ collapses by Proposition 2.7, so we also have
@@
The $\rho$-periodic $\textup{E}_{1}$-page of the real kq-resolution is isomorphic to the $\textup{E}_{1}$-page of the complex kq-resolution.
We can also prove analogous statements using the Brown–Gitler comodules.
@@
$\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k)^{\otimes i}/\rho)\cong\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k)^{\otimes i}).$
Proof.
First, note that we have a diagram in Ext coming from the long exact sequences induced by the cofiber sequence
$\textup{kq}\xrightarrow{\rho}\textup{kq}\to\textup{kq}/\rho$
$\cdots$$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}}/\rho)$$\cdots$$\cdots$$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})\oplus\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$$\cdots$$\scriptstyle{\rho}$$\scriptstyle{0}$
This gives an isomorphism in Ext by the same argument as in Proposition 6.2:
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}}/\rho)\cong\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}}).$
By construction, the Brown–Gitler comodules $B_{0}^{\mathbb{R}}(k)$ are $\rho$-torsion free since they are subcomodules of the dual Steenrod algebra. This extends naturally to tensor factors. The analogous diagram of long exact sequences in Ext gives the desired isomorphism. ∎
These isomorphisms will be helpful for determining hidden extensions in the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(k))$ as they allow us to compare with the computations of [17] recalled in Section 5.
@@
obtained from the Betti realization functor of Remark 2.2 by remembering the $C_{2}$-action given by complex conjugation. This functor has been well studied, see for example [25], and exhibits a close relationship between $\mathbb{R}$-motivic and $C_{2}$-equivariant homotopy theory. It was shown by [33] that
$\text{Be}^{C_{2}}(\textup{kq})_{2,\eta}^{\wedge}\simeq(\text{ko}_{C_{2}})_{2,\eta}^{\wedge}.$
@@
$\text{Ext}^{***}_{\mathcal{A}^{\vee}_{C_{2}}}\cong\text{Ext}^{***}_{\mathcal{A}_{\mathbb{R}}^{\vee}}\oplus\text{Ext}^{***}_{NC}.$
This splitting implies that the $\textup{E}_{2}$-page of the $C_{2}$-equivariant Adams spectral sequence computing the ring of cooperations $\pi_{\star}^{C_{2}}(\text{ko}_{C_{2}}\otimes\text{ko}_{C_{2}})$ contains the $\textup{E}_{2}$-page of the $\textbf{mASS}(\textup{kq}\otimes\textup{kq})$ as a summand. This will be investigated further in joint work with Li, Petersen, and Tatum.
@@ 7. The algebraic Atiyah–Hirzebruch spectral sequence
In this section, we compute the groups $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes i})$ for all $i\geq 0$ by a series of algebraic Atiyah–Hirzebruch spectral sequences. Motivated by the presentation of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ given in [24], we organize these computations by coweight modulo 4. Consequently, our results are best displayed in multiple charts which should be considered simultaneously. Throughout, we use the notation of Section 6.
7.1. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$
We first compute $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$ by the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$. This spectral sequence has signature
$\textup{E}_{1}^{s,f,w,a}=\text{Ext}^{s,f,w}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})\otimes_{\mathbb{M}_{2}^{\mathbb{R}}}\mathbb{M}_{2}^{\mathbb{R}}\{[1],[\overline{\xi}_{1}],[\overline{\tau}_{1}]\}\implies\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}^{s,f,w}(B_{0}^{\mathbb{R}}(1)),$
and the differentials take the form
$d_{1}:\text{E}_{r}^{s,f,w,a}\to\text{E}_{r}^{s-1,f+1,w,a-r}.$
@@ 7.1. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$
Note that in terms of coweight, both of these differentials are determined by multiplication by an element in $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ of coweight 0. This implies that we can represent the $d_{1}$ and $d_{2}$ differentials as maps from one Atiyah–Hirzebruch filtration piece to another in the exact same coweight. In particular, as the coweight $cw\equiv 3\,(4)$ portion of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ was shown to be 0 in 6.1, we can ensure that no differentials land in this vanishing region.
Given the complicated nature of this spectral sequence, we only present charts where the differential is nonzero and only include the relevant Atiyah–Hirzebruch filtrations. For example, since the $d_{1}$-differential is trivial on Atiyah–Hirzebruch filtration 0, we omit this filtration from the charts for the $\textup{E}_{1}$-page.
The $\textup{E}_{1}$-page is of the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ is depicted in Figure 7.1, Figure 7.2, and Figure 7.3. The $\textup{E}_{2}$-page of the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ is depicted in Figure 7.4, Figure 7.5, and Figure 7.6. Differentials are blue, linear with respect to the underlying $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure, and preserve motivic weight.
[…]
@@
This differential is witnessed by the Massey product $\langle\rho,h_{0},h_{1}\rangle$. This Massey product is shown to be equal to $\tau h_{1}$ in [24, Thm 8.1] with zero indeterminacy. ∎
In fact, since the differentials are linear over $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ this determines all $d_{3}$-differentials. For example, we have
@@
It is important to note that since $\tau h_{1}$ has coweight 1 in $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$, and since each Atiyah–Hirzebruch filtration is zero in coweight 3, the $d_{3}$-differential is trivial on the coweight $cw\equiv 2\,(4)$ portion of Atiyah–Hirzebruch filtration 3. This coweight jump requires us to pay close attention to the $d_{3}$-differential. For example, in the coweight $cw\equiv 0\,(4)$ piece of Atiyah–Hirzebruch filtration 3, there is an infinite $\rho$-tower stemming from $\rho[3]$. However, while
@@
Figure 7.7 depicts the $\textup{E}_{3}$-page where the coweight of Atiyah–Hirzebruch filtration 3 is $cw\equiv 0\,(4)$ and the coweight of Atiyah–Hirzebruch filtration 0 is $cw\equiv 1\,(4)$. Figure 7.8 depicts the $\textup{E}_{3}$-page where the coweight of Atiyah–Hirzebruch filtration 3 is $cw\equiv 1\,(4)$ and the coweight of Atiyah–Hirzebruch filtration 0 is $cw\equiv 2\,(4)$.
[…]
Recall from Proposition 4.4 that $\textup{E}_{4}=\textup{E}_{\infty}$ for Atiyah–Hirzebruch filtration reasons. Hidden extensions can be realized using complex Betti realization, linearity over $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$, base change $\otimes\mathbb{C}:\text{SH}(\mathbb{C})\to\text{SH}(\mathbb{R})$, and the isomorphism from Proposition 6.5:
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)/\rho)\cong\text{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)).$
The following result will be useful when depicting the $\textup{E}_{\infty}$-page in charts.